(3,5);x+8y=9
Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b. This is my problem. I have serious difficulties trying to solve these types of equations. This is for Algebra 1 in Axia College online degree program.

Answer by nerdybill(7384)

(3,5);x+8y=9
.
First, we find the slope of:
x+8y=9
Do this by putting it into the "slope-intercept" form:
y = mx +b
where
m is the slope
.
x+8y=9
8y= -x+9
y = (-1/8)x + 9/8
From above, we know our new line has
slope of -1/8
and it crosses:
(3,5)
.
Plug the above data into the "point-slope" form:
y - y1 = m(x - x1)
y - 5 = (-1/8)(x - 3)
y - 5 = (-1/8)x + 3/8
y = (-1/8)x + 3/8 + 5
y = (-1/8)x + 3/8 + 40/8
y = (-1/8)x + 43/8 (this is what they're looking for)